\begin{thebibliography}{10}

\bibitem{BANNERTGROCHENIGSTOCKLER}
Severin Bannert, Karlheinz Grochenig, and Joachim Stockler.
\newblock Discretized {G}abor frames of totally positive functions.
\newblock {\em IEEE Trans. Inform. Theory}, 60(1):159--169, 2014.

\bibitem{BOLCSKEIHLAWATSCH}
H.~Bolcskei and F~Hlawatsch.
\newblock Discrete {Z}ak transforms, polyphase transforms, and applications.
\newblock {\em IEEE Trans. Signal Process.}, 45(4):851--866, 1997.

\bibitem{BORICHEVGROCHENIGLYUBARSKII}
A.~Borichev, K.~Grochenig, and Yu. Lyubarskii.
\newblock Frame constants of {G}abor frames near the critical density.
\newblock {\em J. Math. Pures Appl. (9)}, 94(2):170--182, 2010.

\bibitem{CASAZZA}
P.~G. Casazza, S.~J. Dilworth, E.~Odell, Th. Schlumprecht, and A.~Zsak.
\newblock Coefficient quantization for frames in {B}anach spaces.
\newblock {\em J. Math. Anal. Appl.}, 348(1):66--86, 2008.

\bibitem{CASAZZAHANLARSON}
Peter~G. Casazza, Deguang Han, and David~R. Larson.
\newblock Frames for {B}anach spaces.
\newblock In {\em The functional and harmonic analysis of wavelets and frames
  ({S}an {A}ntonio, {TX}, 1999)}, volume 247 of {\em Contemp. Math.}, pages
  149--182. Amer. Math. Soc., Providence, RI, 1999.

\bibitem{CASAZZAKALTON}
Peter~G. Casazza and Nigel~J. Kalton.
\newblock Roots of complex polynomials and {W}eyl-{H}eisenberg frame sets.
\newblock {\em Proc. Amer. Math. Soc.}, 130(8):2313--2318, 2002.

\bibitem{OLEBOOK}
Ole Christensen.
\newblock {\em Frames and bases: An introductory course}.
\newblock Applied and Numerical Harmonic Analysis. Birkh\"{a}user Boston, Inc.,
  Boston, MA, 2008.

\bibitem{CVETKOVICVETTERLI2}
Zoran Cvetkovic and Martin Vetterli.
\newblock Oversampled filter banks.
\newblock {\em IEEE Trans. Signal Process.}, 46(5):1245--1255, 1998.

\bibitem{CVETKOVICVETTERLI}
Zoran Cvetkovic and Martin Vetterli.
\newblock Tight {W}eyl-{H}eisenberg frames in {$l^2({\bf Z})$}.
\newblock {\em IEEE Trans. Signal Process.}, 46(5):1256--1259, 1998.

\bibitem{DAISUNBOOK}
Xin-Rong Dai and Qiyu Sun.
\newblock The {$abc$}-problem for {G}abor systems and uniform sampling in
  shift-invariant spaces.
\newblock In {\em Excursions in harmonic analysis. {V}ol. 3}, Appl. Numer.
  Harmon. Anal., pages 177--194. Birkh\"{a}user/Springer, Cham, 2015.

\bibitem{DAISUN}
Xin-Rong Dai and Qiyu Sun.
\newblock The {$abc$}-problem for {G}abor systems.
\newblock {\em Mem. Amer. Math. Soc.}, 244(1152):ix+99, 2016.

\bibitem{DAUBECHIES1990}
Ingrid Daubechies.
\newblock The wavelet transform, time-frequency localization and signal
  analysis.
\newblock {\em IEEE Trans. Inform. Theory}, 36(5):961--1005, 1990.

\bibitem{DAUBECHIESBOOK}
Ingrid Daubechies.
\newblock {\em Ten lectures on wavelets}, volume~61 of {\em CBMS-NSF Regional
  Conference Series in Applied Mathematics}.
\newblock Society for Industrial and Applied Mathematics (SIAM), Philadelphia,
  PA, 1992.

\bibitem{DAUBECHIESGROSSMANN}
Ingrid Daubechies and A.~Grossmann.
\newblock Frames in the {B}argmann space of entire functions.
\newblock {\em Comm. Pure Appl. Math.}, 41(2):151--164, 1988.

\bibitem{DAUBECHIESGROSSMANNMEYER}
Ingrid Daubechies, A.~Grossmann, and Y.~Meyer.
\newblock Painless nonorthogonal expansions.
\newblock {\em J. Math. Phys.}, 27(5):1271--1283, 1986.

\bibitem{DUFFIN}
R.~J. Duffin and A.~C. Schaeffer.
\newblock A class of nonharmonic {F}ourier series.
\newblock {\em Trans. Amer. Math. Soc.}, 72:341--366, 1952.

\bibitem{EISNERFREEMAN}
Joseph Eisner and Daniel Freeman.
\newblock Continuous {S}chauder frames for {B}anach spaces.
\newblock {\em J. Fourier Anal. Appl.}, 26(4):Paper No. 66, 30, 2020.

\bibitem{FEICHTINGERKAIBLINGER}
Hans~G. Feichtinger and Norbert Kaiblinger.
\newblock Varying the time-frequency lattice of {G}abor frames.
\newblock {\em Trans. Amer. Math. Soc.}, 356(5):2001--2023, 2004.

\bibitem{FEICHTINGERSTROHMER}
Hans~G. Feichtinger and Thomas Strohmer, editors.
\newblock {\em Gabor analysis and algorithms: Theory and applications}.
\newblock Applied and Numerical Harmonic Analysis. Birkh\"{a}user Boston, Inc.,
  Boston, MA, 1998.

\bibitem{FEICHTINGERSTROHMER2}
Hans~G. Feichtinger and Thomas Strohmer, editors.
\newblock {\em Advances in {G}abor analysis}.
\newblock Applied and Numerical Harmonic Analysis. Birkh\"{a}user Boston, Inc.,
  Boston, MA, 2003.

\bibitem{FLANDRIN}
Patrick Flandrin.
\newblock {\em Time-frequency/time-scale analysis}, volume~10 of {\em Wavelet
  Analysis and its Applications}.
\newblock Academic Press, Inc., San Diego, CA, 1999.

\bibitem{FREEMANODELL}
D.~Freeman, E.~Odell, Th. Schlumprecht, and A.~Zs\'{a}k.
\newblock Unconditional structures of translates for {$L_p(\mathbb{R}^d)$}.
\newblock {\em Israel J. Math.}, 203(1):189--209, 2014.

\bibitem{GOHRONSHEN}
Say~Song Goh, Amos Ron, and Zuowei Shen, editors.
\newblock {\em Gabor and wavelet frames}, volume~10 of {\em Lecture Notes
  Series. Institute for Mathematical Sciences. National University of
  Singapore}.
\newblock World Scientific Publishing Co. Pte. Ltd., Hackensack, NJ, 2007.

\bibitem{GROCHENIG2001}
Karlheinz Grochenig.
\newblock {\em Foundations of time-frequency analysis}.
\newblock Applied and Numerical Harmonic Analysis. Birkh\"{a}user Boston, Inc.,
  Boston, MA, 2001.

\bibitem{GROCHENIGMYSTERY}
Karlheinz Grochenig.
\newblock The mystery of {G}abor frames.
\newblock {\em J. Fourier Anal. Appl.}, 20(4):865--895, 2014.

\bibitem{GROCHENIGROMEROSTOCKLER}
Karlheinz Grochenig, Jose~Luis Romero, and Joachim Stockler.
\newblock Sampling theorems for shift-invariant spaces, {G}abor frames, and
  totally positive functions.
\newblock {\em Invent. Math.}, 211(3):1119--1148, 2018.

\bibitem{GROCHENIGSTOCKLER}
Karlheinz Grochenig and Joachim Stockler.
\newblock Gabor frames and totally positive functions.
\newblock {\em Duke Math. J.}, 162(6):1003--1031, 2013.

\bibitem{GUHAN}
Qing Gu and Deguang Han.
\newblock When a characteristic function generates a {G}abor frame.
\newblock {\em Appl. Comput. Harmon. Anal.}, 24(3):290--309, 2008.

\bibitem{HANMEMOIRS}
Deguang Han and David~R. Larson.
\newblock Frames, bases and group representations.
\newblock {\em Mem. Amer. Math. Soc.}, 147(697):x+94, 2000.

\bibitem{HELAU}
Xing-Gang He and Ka-Sing Lau.
\newblock On the {W}eyl-{H}eisenberg frames generated by simple functions.
\newblock {\em J. Funct. Anal.}, 261(4):1010--1027, 2011.

\bibitem{HEIL2007}
Christopher Heil.
\newblock History and evolution of the density theorem for {G}abor frames.
\newblock {\em J. Fourier Anal. Appl.}, 13(2):113--166, 2007.

\bibitem{JANSSEN1981}
A.~J. E.~M. Janssen.
\newblock Gabor representation of generalized functions.
\newblock {\em J. Math. Anal. Appl.}, 83(2):377--394, 1981.

\bibitem{JANSSEN1982}
A.~J. E.~M. Janssen.
\newblock Bargmann transform, {Z}ak transform, and coherent states.
\newblock {\em J. Math. Phys.}, 23(5):720--731, 1982.

\bibitem{JANSSEN1994}
A.~J. E.~M. Janssen.
\newblock Signal analytic proofs of two basic results on lattice expansions.
\newblock {\em Appl. Comput. Harmon. Anal.}, 1(4):350--354, 1994.

\bibitem{JANSSENAEJM}
A.~J. E.~M. Janssen.
\newblock Signal analytic proofs of two basic results on lattice expansions.
\newblock {\em Appl. Comput. Harmon. Anal.}, 1(4):350--354, 1994.

\bibitem{JANSSEN1995}
A.~J. E.~M. Janssen.
\newblock Duality and biorthogonality for {W}eyl-{H}eisenberg frames.
\newblock {\em J. Fourier Anal. Appl.}, 1(4):403--436, 1995.

\bibitem{JANSSEN1996}
A.~J. E.~M. Janssen.
\newblock Some {W}eyl-{H}eisenberg frame bound calculations.
\newblock {\em Indag. Math. (N.S.)}, 7(2):165--183, 1996.

\bibitem{JANSSEN2000}
A.~J. E.~M. Janssen.
\newblock Representations of {G}abor frame operators.
\newblock In {\em Twentieth century harmonic analysis---a celebration ({I}l
  {C}iocco, 2000)}, volume~33 of {\em NATO Sci. Ser. II Math. Phys. Chem.},
  pages 73--101. Kluwer Acad. Publ., Dordrecht, 2001.

\bibitem{JANSSEN2003}
A.~J. E.~M. Janssen.
\newblock On generating tight {G}abor frames at critical density.
\newblock {\em J. Fourier Anal. Appl.}, 9(2):175--214, 2003.

\bibitem{JANSSENSTROHMER2002}
A.~J. E.~M. Janssen and Thomas Strohmer.
\newblock Hyperbolic secants yield {G}abor frames.
\newblock {\em Appl. Comput. Harmon. Anal.}, 12(2):259--267, 2002.

\bibitem{KLOOSSTOCKLER}
Tobias Kloos and Joachim Stockler.
\newblock Zak transforms and {G}abor frames of totally positive functions and
  exponential {B}-splines.
\newblock {\em J. Approx. Theory}, 184:209--237, 2014.

\bibitem{KRISHNA2}
K.~Mahesh Krishna.
\newblock Discrete and continuous {W}elch bounds for {B}anach spaces with
  applications.
\newblock {\em arXiv:2201.00980v1 [math.FA] 4 January}, 2022.

\bibitem{KRISHNA}
K.~Mahesh Krishna.
\newblock Feichtinger conjectures, {$R_\varepsilon$}-conjectures and
  {W}eaver’s conjectures for {B}anach spaces.
\newblock {\em arXiv:2201.00125v1 [math.FA] 1 January}, 2022.

\bibitem{KRISHNAJOHNSON}
K.~Mahesh Krishna and P.~Sam Johnson.
\newblock Towards characterizations of approximate {S}chauder frame and its
  duals for {B}anach spaces.
\newblock {\em J. Pseudo-Differ. Oper. Appl.}, 12(1):Paper No. 9, 13, 2021.

\bibitem{KRISHNAJOHNSON2}
K.~Mahesh Krishna and Johnson~P. Sam.
\newblock Dilation theorem for p-approximate {S}chauder frames for separable
  {B}anach spaces.
\newblock {\em arXiv:2011.02188v1 [math.FA] 23 November}, 2020.

\bibitem{KRISHNAJOHNSON3}
K.~Mahesh Krishna and Johnson~P. Sam.
\newblock Approximately dual p-approximate {S}chauder frames.
\newblock {\em arXiv:2110.10121v1 [math.FA] 19 October}, 2021.

\bibitem{LANDAU}
H.~Landau.
\newblock On the density of phase space expansions.
\newblock {\em IEEE Trans. Inform. Theory}, 39(5):1152--1156, 1993.

\bibitem{LEMVIG2017}
Jakob Lemvig.
\newblock On some {H}ermite series identities and their applications to {G}abor
  analysis.
\newblock {\em Monatsh. Math.}, 182(4):899--912, 2017.

\bibitem{LEMVIGHAAHR}
Jakob Lemvig and Kamilla Haahr~Nielsen.
\newblock Counterexamples to the {B}-spline conjecture for {G}abor frames.
\newblock {\em J. Fourier Anal. Appl.}, 22(6):1440--1451, 2016.

\bibitem{LOPEZHAN}
Jerry Lopez and Deguang Han.
\newblock Discrete {G}abor frames in {$\ell^2(\Bbb{Z}^d)$}.
\newblock {\em Proc. Amer. Math. Soc.}, 141(11):3839--3851, 2013.

\bibitem{LYUBARSKIINES}
Yurii Lyubarskii and Preben Gr\aa~berg Nes.
\newblock Gabor frames with rational density.
\newblock {\em Appl. Comput. Harmon. Anal.}, 34(3):488--494, 2013.

\bibitem{LYUBARSKIISEIP}
Yurii~I. Lyubarskii and Kristian Seip.
\newblock Convergence and summability of {G}abor expansions at the {N}yquist
  density.
\newblock {\em J. Fourier Anal. Appl.}, 5(2-3):127--157, 1999.

\bibitem{LYUBARSKII}
Yu.~I. Lyubarski\u{\i}.
\newblock Frames in the {B}argmann space of entire functions.
\newblock In {\em Entire and subharmonic functions}, volume~11 of {\em Adv.
  Soviet Math.}, pages 167--180. Amer. Math. Soc., Providence, RI, 1992.

\bibitem{MORRISLU}
J.~M. Morris and Y.~Lu.
\newblock Generalized {G}abor expansions of discrete-time signals in {$l^2({\bf
  Z})$} via biorthogonal-like sequences.
\newblock {\em IEEE Trans. Signal Process.}, 44(6):1378--1391, 1996.

\bibitem{OKOUDJOU}
Kasso~A. Okoudjou.
\newblock An invitation to {G}abor analysis.
\newblock {\em Notices Amer. Math. Soc.}, 66(6):808--819, 2019.

\bibitem{RONSHEN}
Amos Ron and Zuowei Shen.
\newblock Weyl-{H}eisenberg frames and {R}iesz bases in {$L_2(\bold R^d)$}.
\newblock {\em Duke Math. J.}, 89(2):237--282, 1997.

\bibitem{SEIP1992}
Kristian Seip.
\newblock Density theorems for sampling and interpolation in the
  {B}argmann-{F}ock space. {I}.
\newblock {\em J. Reine Angew. Math.}, 429:91--106, 1992.

\bibitem{SEIPWALLSTEN1992}
Kristian Seip and Robert Wallsten.
\newblock Density theorems for sampling and interpolation in the
  {B}argmann-{F}ock space. {II}.
\newblock {\em J. Reine Angew. Math.}, 429:107--113, 1992.

\bibitem{THOMAS}
S.~M. Thomas.
\newblock Approximate {S}chauder frames for {$\mathbb{R}^n$, Masters Thesis,
  St. Louis University, St. Louis, MO}.
\newblock 2012.

\end{thebibliography}

\documentclass{amsart}
\usepackage{setspace}
\usepackage{a4}
\usepackage{amssymb,amsmath,amsthm,latexsym}
\usepackage{amsfonts}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{textcomp}
\usepackage{cite}
\usepackage{enumerate}
\usepackage[mathscr]{euscript}
\usepackage{mathtools}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{definition}[theorem]{Definition}

\newtheorem{problem}[theorem]{Problem}

\title{This is the title}
\usepackage{amssymb}
\usepackage{amssymb}
\usepackage{amssymb}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{hyperref}
\usepackage{enumerate}
\usepackage{mathtools}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{hyperref}
\usepackage{amssymb}
\usepackage{amssymb}
\usepackage{amssymb}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{hyperref}
\usepackage{enumerate}
\usepackage{mathtools}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{graphicx}
\usepackage{textcomp}
\usetikzlibrary{cd}
\usepackage{fancyhdr}

\begin{document}
\begin{center}
{\bf{THE $(abc,pqr)$-PROBLEM FOR APPROXIMATE SCHAUDER FRAMES FOR  BANACH SPACES}}\\
\textbf{K. MAHESH KRISHNA} \\
Post Doctoral Fellow \\
Statistics and Mathematics Unit\\
Indian Statistical Institute, Bangalore Centre\\
Karnataka 560 059 India\\
Email: kmaheshak@gmail.com \\
Date: \today
\end{center}
\hrule
\vspace{0.5cm}
\textbf{Abstract}: Motivated from the complete solution of important $abc$-problem for Gabor system for the Hilbert space $\mathcal{L}^2(\mathbb{R})$ by Dai and Sun [\textit{Memoirs of Amer. Math. Soc., 2016}] and from the existential result of approximate Schauder frames for $\mathcal{L}^p(\mathbb{R})$ using translation operators on $\mathcal{L}^p(\mathbb{R})$ by Freeman, Odell, Schlumprecht, and Zsak [\textit{Israel. J. Math, 2014}], we formulate $(abc, pqr)$-problem for approximate Schauder frames for Banach spaces  $\mathcal{L}^p(\mathbb{R})$, $1<p<\infty$.

\textbf{Keywords}: $abc$-problem for Gabor system,  Approximate Schauder Frame.

\textbf{Mathematics Subject Classification (2020)}: 42C15, 46E30.

\tableofcontents 

\section{Introduction}
Let   $\mathcal{H}$ be a  Hilbert space over $\mathbb{C}$. Recall that \cite{DUFFIN, OLEBOOK, HANMEMOIRS}
a sequence $\{\tau_n\}_n$ in  $\mathcal{H}$ is said to be a
frame for $\mathcal{H}$ if there exist $a,b>0$ such that 
\begin{align*}
a\|h\|^2 \leq \sum_{n=1}^\infty |\langle h, \tau_n\rangle|^2\leq b\|h\|^2, \quad \forall h \in \mathcal{H}.
\end{align*}
Among various examples of frames for Hilbert spaces, one of the most studied classes of frames for the Hilbert space $\mathcal{L}^2(\mathbb{R})$ is the frames arising from the Gabor systems 
\begin{align*}
	\{E_{mb}T_{na}g\}_{m,n \in \mathbb{Z}},
\end{align*}
where $g\in \mathcal{L}^2(\mathbb{R})$ and for $c, d \in \mathbb{R}$,
\begin{align*}
&T_c:	\mathcal{L}^2(\mathbb{R})\ni f \mapsto T_cf \in \mathcal{L}^2(\mathbb{R});~  T_cf: \mathbb{R} \ni x \mapsto ( T_cf)(x)\coloneqq f(x-c)\in \mathbb{C},\\
&E_d:	\mathcal{L}^2(\mathbb{R})\ni f \mapsto E_df \in \mathcal{L}^2(\mathbb{R});~  E_df: \mathbb{R} \ni x \mapsto ( E_df)(x)\coloneqq e^{2\pi i d x}f(x)\in \mathbb{C}.
\end{align*}
A detailed analysis of Gabor frames for Hilbert spaces can be found in \cite{FLANDRIN, DAUBECHIESBOOK, FEICHTINGERSTROHMER, FEICHTINGERSTROHMER2, GOHRONSHEN, GROCHENIGMYSTERY, OKOUDJOU}.   A fundamental problem in Gabor analysis is to characterize parameters $a>0$, $b>0$ and the function $g$ such that $	\{E_{mb}T_{na}g\}_{m,n \in \mathbb{Z}}$ is a frame for $\mathcal{L}^2(\mathbb{R})$. The problem has been solved  fully or partially for certain classes of functions in \cite{JANSSEN1981, JANSSEN1982, FEICHTINGERKAIBLINGER, GROCHENIG2001, JANSSEN2000, DAUBECHIESGROSSMANNMEYER, GROCHENIGSTOCKLER, GUHAN, RONSHEN, JANSSENSTROHMER2002, JANSSEN1996, JANSSEN2003, LYUBARSKII, SEIPWALLSTEN1992, SEIP1992, LYUBARSKIISEIP, BORICHEVGROCHENIGLYUBARSKII, JANSSEN1994, DAUBECHIESGROSSMANN, GROCHENIGROMEROSTOCKLER, LYUBARSKIINES, LEMVIG2017, LEMVIGHAAHR, GUHAN, HELAU, CASAZZAKALTON, HEIL2007, JANSSEN1995, JANSSENAEJM, DAUBECHIES1990, LANDAU, KLOOSSTOCKLER}. Let $c>0$. When $g=\chi_{[0,c]}$, the characteristic function on $[0,c]$, the problem of characterizing  parameters $a>0$, $b>0$ and $c>0$ such that 
\begin{align*}
	\{E_{mb}T_{na}\chi_{[0,c]}\}_{m,n \in \mathbb{Z}}
\end{align*}
is a frame for $\mathcal{L}^2(\mathbb{R})$ is known as $abc$-problem for Gabor system. A list of partial answers for this are obtained in \cite{LEMVIGHAAHR, GROCHENIGMYSTERY, LEMVIG2017, LYUBARSKIINES, GUHAN, HELAU, HEIL2007, CASAZZAKALTON, FEICHTINGERSTROHMER2}. In 2013, Dai and Sun fully resolved this problem in their Memoirs  
\cite{DAISUN, DAISUNBOOK}. In this paper, we formulate  a similar problem for approximate Schauder frames for function spaces.

\section{The $(abc, pqr)$-Problem for Approximate Schauder Frames}
We are motivated from the following two results.
\begin{enumerate}[\upshape(i)]
	\item The complete solution of $abc$-problem for Gabor system by Dai and Sun \cite{DAISUN}.
	\item For every $2<p<\infty$, there is an  approximate Schauder frame for $\mathcal{L}^p(\mathbb{R})$ obtained  using translation operators on $\mathcal{L}^p(\mathbb{R})$ \cite{FREEMANODELL}.
\end{enumerate}
  Generalized Fourier expansion which resulted from the theory of frames for Hilbert spaces led the notion of framing in \cite{CASAZZAHANLARSON} which is generalized to Schauder frames in \cite{CASAZZA}   which is further generalized in \cite{FREEMANODELL, THOMAS} to approximate Schauder frames defined as follows.
\begin{definition}\cite{FREEMANODELL, THOMAS}\label{ASFDEF}
 Let $\mathcal{X}$ be a separable Banach space and $\mathcal{X}^*$ be its dual.	Let $\{\tau_n\}_n$ be a sequence in  $\mathcal{X}$ and 	$\{f_n\}_n$ be a sequence in  $\mathcal{X}^*.$ The pair $ (\{f_n \}_{n}, \{\tau_n \}_{n}) $ is said to be an \textbf{approximate Schauder frame (ASF)} for $\mathcal{X}$ if the map 
	\begin{align*}
	S_{f, \tau}:\mathcal{X}\ni x \mapsto S_{f, \tau}x\coloneqq \sum_{n=1}^\infty
	f_n(x)\tau_n \in
	\mathcal{X}
	\end{align*}
	is a well-defined bounded linear, invertible operator.
\end{definition} 
We now consider the Banach space $\mathcal{X}=\mathcal{L}^p(\mathbb{R})$, for $1<p<\infty$. Let $q$ be the conjugate index of $p$. Since dual of $\mathcal{L}^p(\mathbb{R})$ is isometrically isomorphic to $\mathcal{L}^q(\mathbb{R})$, given $\phi \in \mathcal{X}^*$, there exists a unique $\omega\in \mathcal{L}^q(\mathbb{R})$ such that 
\begin{align*}
\phi(u)=\int_\mathbb{R}u(\alpha)\omega(\alpha)\,d\alpha \eqqcolon [u,\omega], \quad \forall u \in \mathcal{L}^p(\mathbb{R}).
\end{align*}

  In particular, if $\{f_n\}_n$ is a sequence in  $\mathcal{X}^*$, then there exists a unique sequence $\{\omega_n\}_n$ in  $\mathcal{X}$ such that 
  \begin{align*}
  	f_n(x)=[x, \omega_n], \quad \forall x \in \mathcal{X}, \forall n \in \mathbb{N}.
  \end{align*}
 Thus asking whether a pair $ (\{f_n \}_{n}, \{\tau_n \}_{n}) $ is  an ASF for $\mathcal{X}$ is same as asking whether the map 
 	\begin{align*}
 	S_{\omega, \tau}:\mathcal{X}\ni x \mapsto S_{\omega, \tau}x\coloneqq \sum_{n=1}^\infty
 [x, \omega_n]	\tau_n \in
 	\mathcal{X}
 \end{align*}
 is a well-defined bounded linear, invertible operator. We then have the following problem.
   \begin{problem}\label{GABORBANACH}
 \textbf{Let $1<p<\infty$ and $q$ be the conjugate index of $p$. Characterize parameters $a>0$, 	$b>0$, $p>0$, $q>0$ and functions $g\in \mathcal{L}^p(\mathbb{R})$, $h\in \mathcal{L}^q(\mathbb{R})$ such that the pair 
  \begin{align*}
  	(\{\tilde{E}_{mq}\tilde{T}_{np}h\}_{m,n \in \mathbb{Z}}, 	\{E_{mb}T_{na}g\}_{m,n \in \mathbb{Z}}),
  \end{align*}
is an ASF for $\mathcal{L}^p(\mathbb{R})$, where 
\begin{align*}
	&T_c:	\mathcal{L}^p(\mathbb{R})\ni f \mapsto T_cf \in \mathcal{L}^p(\mathbb{R});~  T_cf: \mathbb{R} \ni x \mapsto ( T_cf)(x)\coloneqq f(x-c)\in \mathbb{C}, ~ \text{ for } c\in \mathbb{R},\\
	&E_d:	\mathcal{L}^p(\mathbb{R})\ni f \mapsto E_df \in \mathcal{L}^p(\mathbb{R});~  E_df: \mathbb{R} \ni x \mapsto ( E_df)(x)\coloneqq e^{2\pi i d x}f(x)\in \mathbb{C}, ~ \text{ for } d\in \mathbb{R}
\end{align*}
and 
\begin{align*}
	&\tilde{T}_s:	\mathcal{L}^q(\mathbb{R})\ni f \mapsto \tilde{T}_sf \in \mathcal{L}^q(\mathbb{R});~  \tilde{T}_sf: \mathbb{R} \ni x \mapsto ( \tilde{T}_sf)(x)\coloneqq f(x-s)\in \mathbb{C}, ~ \text{ for } s\in \mathbb{R},\\
	&\tilde{E}_t:	\mathcal{L}^q(\mathbb{R})\ni f \mapsto \tilde{E}_tf \in \mathcal{L}^q(\mathbb{R});~  \tilde{E}_tf: \mathbb{R} \ni x \mapsto ( \tilde{E}_tf)(x)\coloneqq e^{2\pi i t x}f(x)\in \mathbb{C}, ~ \text{ for } t\in \mathbb{R}.
\end{align*}}
  \end{problem}
Note that all operators $T_c$, $E_c$,  $\tilde{T}_s $,  $\tilde{E}_t $ are invertible isometries. We also note that Problem \ref{GABORBANACH} can be stated for space  $\mathcal{L}^p[a,b].$ A particular case of Problem  \ref{GABORBANACH}  is the following $(abc, pqr)$-problem.
  \begin{problem}\textbf{(The $(abc, pqr)$-problem for approximate Schauder frames)}\label{ABCPQR}
   \textbf{Let $1<p<\infty$ and $q$ be the conjugate index of $p$. Characterize parameters $a>0$, 	$b>0$, $c>0$, $p>0$, $q>0$, $r>0$   such that the pair 
  	\begin{align*}
  		(\{\tilde{E}_{mq}\tilde{T}_{np}\chi_{[0,r]}\}_{m,n \in \mathbb{Z}}, 	\{E_{mb}T_{na}\chi_{[0,c]}\}_{m,n \in \mathbb{Z}}),
  	\end{align*}
  	is an ASF for $\mathcal{L}^p(\mathbb{R})$}.	
  \end{problem}
 We remark that we can formulate problems similar to Problems \ref{GABORBANACH}, \ref{ABCPQR}  for p-approximate Schauder frames, p-approximate Bessel sequences,  p-approximate orthonormal bases and p-approximate Riesz bases \cite{KRISHNAJOHNSON, KRISHNA, KRISHNAJOHNSON2, KRISHNAJOHNSON3}. Following is one such in the discrete case. Our problem is motivated from the discrete Gabor analysis in $\ell^2(\mathbb{Z})$ \cite{LOPEZHAN, CVETKOVICVETTERLI, MORRISLU, BOLCSKEIHLAWATSCH, CVETKOVICVETTERLI2, BANNERTGROCHENIGSTOCKLER}.
 \begin{problem}\textbf{(The $(MN, PQ)$-problem for p-approximate Schauder frames)}
 \textbf{Let $1<p<\infty$ and $q$ be the conjugate index of $p$. Characterize natural numbers $N,M,P,Q$ and sequences $\{x_j\}_{j\in \mathbb{Z}}\in \ell^p(\mathbb{Z})$, $\{y_j\}_{j\in \mathbb{Z}}\in \ell^q(\mathbb{Z})$ such that the pair 
 	\begin{align*}
 	(\{	\tilde{E}_{m/P}\tilde{T}_{nQ}\{y_j\}_{j\in \mathbb{Z}}\}_{0\leq m\leq P-1, n\in \mathbb{Z}}, \{E_{m/M}T_{nN}\{x_j\}_{j\in \mathbb{Z}}\}_{0\leq m\leq M-1, n\in \mathbb{Z}})
 	\end{align*}
 is a 
 \begin{enumerate}[\upshape(i)]
 	\item p-orthonormal basis for $ \ell^p(\mathbb{Z})$.
 	\item p-approximate Riesz basis for $ \ell^p(\mathbb{Z})$.
 	\item p-approximate Schauder frame for $ \ell^p(\mathbb{Z})$.
 	\item p-approximate Bessel sequence for $ \ell^p(\mathbb{Z})$, 
 \end{enumerate}
where
\begin{align*}
&	\tilde{E}_{m/P}\tilde{T}_{nQ}:  \ell^q(\mathbb{Z})\ni \{a_j\}_{j\in \mathbb{Z}}\mapsto \tilde{E}_{m/P}\tilde{T}_{nQ}\{a_j\}_{j\in \mathbb{Z}}\coloneqq  \{e^{2\pi ijm/P}a_{j-nQ}\}_{j\in \mathbb{Z}}\in  \ell^q(\mathbb{Z}),\\
&	E_{m/M}T_{nN}:  \ell^p(\mathbb{Z})\ni \{b_j\}_{j\in \mathbb{Z}}\mapsto E_{m/M}T_{nN}\{a_j\}_{j\in \mathbb{Z}}\coloneqq  \{e^{2\pi ijm/M}b_{j-nN}\}_{j\in \mathbb{Z}}\in  \ell^p(\mathbb{Z}).
\end{align*}}
 \end{problem} 
 Next we formulate the continuous version of Problem  \ref{GABORBANACH} (see  \cite{KRISHNA2} for the definition of continuous approximate Schauder frame and \cite{EISNERFREEMAN} for the definition of continuous Schauder frame).
  \begin{problem}
   \textbf{Let $1<p<\infty$ and $q$ be the conjugate index of $p$. Characterize  functions $g\in \mathcal{L}^p(\mathbb{R})$, $h\in \mathcal{L}^q(\mathbb{R})$ such that the pair 
  	\begin{align*}
  		(\{	\tilde{E}_b	\tilde{T}_ah\}_{a,b \in \mathbb{R}}, \{	E_b	T_ag\}_{a,b \in \mathbb{R}})
  	\end{align*}
  is a continuous approximate Schauder frame for $\mathcal{L}^p(\mathbb{R})$}.
\end{problem}

 \bibliographystyle{plain}
 \bibliography{reference.bib}

\end{document}